OpenStudy (anonymous):

x/y = 4/5 z/y = 2/5 what is z/y?

hero (hero):

What ? you already have z/y = 2/5....

OpenStudy (anonymous):

sorry i meant x/z.

hero (hero):

Okay....

hero (hero):

wait...it's just 4/2 = 2

hero (hero):

to find x/z...just do x/y *y/z....the y's cancel...

OpenStudy (anonymous):

oh. just making sure. because my cousin said it wasn't. thankyouuu

OpenStudy (anonymous):

ha ha ha... 2/5!! lol

hero (hero):

yeah genius..we already discussed that...He wanted x/z instead...

OpenStudy (anonymous):

i'm a girl but okay. lol.

hero (hero):

OOps

OpenStudy (anonymous):

:)